OpenStudy (anonymous):

3x + 5y = 14 Line that passes through

OpenStudy (anonymous):

do you need this in slope intercept form

OpenStudy (anonymous):

standard form

OpenStudy (anonymous):

like x+y=b

OpenStudy (anonymous):

well its already in standard form

OpenStudy (anonymous):

Oh i guess slope intercept form haha sorry

OpenStudy (anonymous):

then for slope int. you will need the form y=mx+b

OpenStudy (anonymous):

in order to get that we have to get y alone 3x+5y=14 -3x -3x _________________________________ 5y=-3x+14 5y/y= -3x/5+14/5 y=? Do you know how solve from here